# Learning with less: label-efficient land cover classification at very high spatial resolution using self-supervised deep learning

### v1

@@ Abstract
Deep learning semantic segmentation methods have shown promising performance for very high resolution land cover classification, but the challenge of collecting large volumes of representative training data creates a significant barrier to widespread adoption of such models for meter-scale land cover mapping over large areas. In this study, we present a novel label-efficient approach for statewide land cover classification using only 1,000 annotated reference image patches with self-supervised deep learning. We use the “Bootstrap Your Own Latent” pre-training strategy with a large amount of unlabeled color-infrared aerial images (377,921 256256 pixel patches) to pre-train a ResNet-101 convolutional encoder. The learned encoder weights were subsequently transferred into multiple deep semantic segmentation architectures (FCN, U-Net, Attention U-Net, DeepLabV3+, UPerNet, PAN), which were then fine-tuned using very small training dataset sizes with cross-validation (250, 500, 750 patches). Among the fine-tuned models, we obtained the 87.14% overall accuracy and 75.58% macro F1 score using an ensemble of the best performing U-Net models for comprehensive , 8-class land cover mapping, covering more than 123 billion pixels over the state of Mississippi, USA. Detailed qualitative and quantitative analysis revealed accurate mapping of open water and forested areas, while highlighting challenges in accurate delineation between cropland, herbaceous, and barren land cover types. These results show that self-supervised learning is an effective strategy for reducing the need for large volumes of manually annotated data, directly addressing a major limitation to high spatial resolution land cover mapping at scale.
@@ 1 Introduction
Land cover data products provide valuable information about the characteristics of the Earth’s surface for a wide variety of applications, including agriculture, urban planning, and environmental monitoring [1, 2, 3, 4, 5]. Research into automated methods for land cover classification from remotely sensed data has been ongoing for several decades, with a majority of methods relying on supervised learning techniques to classify land cover using multispectral aerial or satellite imagery [6, 7, 8, 9, 10]. Historically, pixel-wise and object-based image analysis (OBIA) classification methods have been the predominant approach for land cover classification, where spatial units (i.e., pixels or objects) in the image are classified based on their spectral reflectance, hand-crafted features, and ancillary data from other geospatial datasets [11, 12]. Very high spatial resolution (VHSR, ) land cover products offer detailed information about the Earth’s surface with high spatial accuracy, which is valuable for applications that require fine-grained information about land cover types and their spatial distribution [13, 14]. However, the creation of such VHSR data products is inherently challenging due to the complexity of the land cover patterns at these scales and the increased spectral variability within classes [15, 16, 17]. In this context, the application of deep convolutional neural networks (CNNs) to land cover classification tasks has gained significant attention in recent years, as these models excel at extracting spatial and spectral features from imagery in a data-driven manner, making them well-suited for VHSR land cover classification tasks [18, 19, 20]. Nevertheless, these models require far more labeled training data compared to traditional machine learning methods, serving as a significant barrier to their widespread adoption for operational land cover mapping at VHSR [21, 22].
@@
This need for large amounts of ground-truth data for accurate training of deep learning models has led to extensive research into methods for reducing the amount of data necessary to train these models [23, 24, 25]. Most of these methods leverage transfer learning procedures, where the parameters of a model pre-trained using a large dataset are used to initialize a model that is subsequently fine-tuned on a smaller dataset that is more relevant to the target task [26, 27]. For example, a common approach to transfer learning for land cover classification tasks is to pre-train a model on a large general-purpose image classification dataset (e.g., ImageNet [28]) and then fine-tune the model on a smaller dataset of interest [29, 30, 31, 32]. Despite being effective, the domain shift between the source and target datasets can hinder the model’s performance, and the extent of this domain shift between general-purpose image classification datasets and VHSR remote sensing datasets is not well understood [33]. As fully-supervised pre-training requires large-scale labeled datasets, which are relatively scarce in the remote sensing domain for VHSR imagery, there is a need for alternative pre-training strategies that reduce the reliance on labeled data [34, 35].
@@
More recently, self-supervised pre-training methods have been proposed as an alternative to traditional fully-supervised pre-training methods, enabling models to learn useful spatial and spectral information from input imagery through the use of pretext tasks that do not require labeled data [36, 37, 38]. These techniques are particularly attractive for remote sensing applications, as large amounts of labeled imagery datasets are often unavailable, but source imagery is widely available due to the growing number of Earth observation satellites and aerial imagery platforms [39]. However, few studies have rigorously investigated the applicability of these methods to operational land cover mapping at VHSR, especially in scenarios where ground-truth data are scarce. Most research evaluating self-supervised models pre-trained with remote sensing data uses benchmark VHSR land cover datasets as a proxy to evaluate the quality of pre-trained models, with the ISPRS Potsdam [40] and Vaihingen [41] datasets seeing wide use for this purpose [42, 43, 44, 45]. However, these datasets contain relatively large amounts of labeled training data, and thus may not necessarily be representative of real-world scenarios where no available VHSR land cover data exists and costly manual annotation is required to create such data (e.g., by experience, a single 256256 image patch can take upwards of 30 minutes to label well). Furthermore, while there is growing momentum in the development of large remote sensing foundation models using self-supervised techniques [46, 39], there is still a critical need to systematically assess how these approaches can enhance VHSR land cover mapping performance in practical contexts, a gap this study directly addresses.
@@
Self-supervised approaches inspired by masked language modeling in natural language processing have been shown to be strong approaches for learning useful feature representations from images without the need for labeled data [47, 48]. Masked image modeling (MIM) methods take advantage of the Vision Transformer (ViT) patching mechanism to randomly mask out a portion of the input image and then train the model to reconstruct the missing patches from the visible patches [49, 50]. MIM techniques have become the predominant self-supervised pre-training paradigm for creating remote sensing foundation models, which are large pre-trained models that can be adapted for a wide variety of tasks and domains that ideally outperform models trained from scratch or use more traditional transfer learning approaches [51, 45]. However, while powerful, ViTs lack the inductive biases toward imagery that make CNNs effective at extracting spatial features from images [52]. This is problematic, as ViTs typically only begin to outperform CNNs when pre-trained on very large datasets [53], a practical consideration for researchers and practitioners who must balance both computational resources and data availability when developing models. Self-distillation techniques, such as Bootstrap Your Own Latent (BYOL) [54] and DINO [55], are alternatives to MIM pre-training that can be used to learn robust and transferable feature representations without the need for negative samples, large batch sizes, or ViT backbones, which are hard requirements for many other self-supervised learning methods [38, 56]. Such approaches largely revolve around maximizing the similarity between feature representations of different augmented views of the same input using a teacher-student framework [57]. While the student network is trained directly using standard backpropagation and gradient descent algorithms, the teacher’s parameters are updated using a momentum-based rule derived from the student’s parameters. Self-distillation methods are attractive for remote sensing applications due to their inherent flexibility and ability to learn strong feature representations using a variety of backbone architectures, including CNNs [58]. Given the architectural constraints of MIM and the computational demands of contrastive methods, self-distillation techniques offer a compelling, architecture-agnostic solution, suitable for pre-training CNNs on domain-specific, moderate-sized datasets [59].
@@
In this study, we propose an innovative label-efficient framework that addresses the challenge of operational land cover mapping at resolution under label-scarce scenarios through self-supervised learning. Specifically, our framework used BYOL as a pretext task during pre-training, enabling a ResNet-101 convolutional encoder network to learn to extract deep features from unlabeled 377,921 256256 patches of color-infrared aerial imagery without the need for full supervision using labeled data. We then transferred the learned encoder weights into multiple semantic segmentation architectures (FCN, U-Net, Attention U-Net, DeepLabV3+, UPerNet, PAN) that are subsequently fine-tuned using very small training dataset sizes (250, 500, and 750 256256 patches) using cross-validation. We classified 8 classes of land cover (open water, impervious structures, impervious surfaces, barren land, forest/woody vegetation, herbaceous/low vegetation, cultivated crops, and unclassified (i.e., shadow)) over the entire state of Mississippi, USA using an ensemble of the best-performing models, totaling over 123 billion pixels. Validation and accuracy assessment were performed using a spatially independent test set of 25,000 randomly sampled points with manually verified labels to ensure robust and unbiased evaluation of our approach and the resulting land cover map. We demonstrated that these approaches can be used to train deep semantic segmentation models that are accurate, generalizable, and capable of yielding reasonable performance even under extreme low-data scenarios, even when only 3 bands (red, green, and near-infrared) are available in the input imagery. Our ultimate aim is to reduce practical barriers to producing high spatial resolution land cover maps while contributing to ongoing efforts to evaluate and contextualize the implications of self-supervised learning techniques for remote sensing purposes.
@@ 2.1 NAIP aerial imagery
The USDA’s National Agricultural Imagery Program (NAIP) is a U.S. federal project that provides a publicly available archive of high-resolution multispectral imagery over the conterminous United States, with a spatial resolution of , a four-band spectral resolution (near-infrared, red, green, and blue), delivered as 8-bit orthorectified tiles [60]. NAIP imagery is collected on a state-by-state basis during the agricultural growing season, and each state is imaged once every 2–3 years. The lack of use restrictions on NAIP imagery has made it a popular choice for high spatial resolution remote sensing applications in the United States, including building footprint extraction [61], forest measurements [62, 63], and land cover classification [64, 65, 66, 67]. Color-infrared (CIR) composites are generated using the near-infrared, red, and green bands to emphasize vegetation features in the imagery. All 2023 CIR composites were obtained as county-level composites from the USDA NRCS Geospatial Data Gateway [68] in MrSID format in the appropriate UTM zone for each county (UTM Zone 15N for Western Mississippi, UTM Zone 16N for Eastern Mississippi) at resolution. Each county-level composite was converted to GeoTiff format, reprojected to the Mississippi Transverse Mercator coordinate reference system (EPSG:3813), and resampled to resolution using bilinear interpolation. The total size of the NAIP imagery used in this work was , comprising 82 county-level GeoTiff files, an average of per county, with individual county rasters ranging in size from to . A visualization of the NAIP imagery used over the study area is provided in Figure 1. We selected three-band images for compatibility with BYOL, which is designed for three-channel inputs, and for easy comparison with models using ImageNet pre-trained weights, which also expect three-channel inputs.
@@ 2.2 NLCD land cover
The stratification approach proposed in this work (Section 3.1) relies on an existing land cover product to select strata from which training samples are drawn. For this purpose, we used the USGS’s National Land Cover Database (NLCD) to provide the necessary land cover information for stratification. The NLCD has historically been among the most widely used land cover datasets in the United States, supporting numerous applications in human and ecosystem health studies, urban planning, natural hazard assessment, hydrology, and land management [69]. The latest version of the NLCD (Annual NLCD) provides a yearly 20-class land cover product primarily derived from Landsat imagery at resolution from 1985 to 2024 [70]. While the spatial resolution of the NLCD land cover product is substantially coarser than the land cover product we aim to produce, the NLCD provides a comprehensive land cover product at a national scale and has a history of consistent use in the literature [71, 72, 73], making it a suitable choice to provide land cover information to guide the stratification process. To align with the collection year of the NAIP imagery used in this work, we use 2023 data over the state of Mississippi.

### v2

@@ Abstract
Deep learning semantic segmentation methods have shown promising performance for very high resolution land cover classification, but the challenge of collecting large volumes of representative training data creates a significant barrier to widespread adoption of such models for meter-scale land cover mapping over large areas. In this study, we present a novel label-efficient approach for statewide land cover classification using only 1,000 annotated reference image patches with self-supervised deep learning. We use the “Bootstrap Your Own Latent” pre-training strategy with a large amount of unlabeled color-infrared aerial images (377,921 patches of 256256 pixels at resolution) to pre-train a ResNet-101 convolutional encoder. The learned encoder weights were subsequently transferred into multiple deep semantic segmentation architectures (FCN, U-Net, Attention U-Net, DeepLabV3+, UPerNet, PAN), which were then fine-tuned using very small training dataset sizes with cross-validation (250, 500, 750 patches). Among the fine-tuned models, we obtained 87.14% overall accuracy and 75.58% macro F1 score using an ensemble of the best-performing U-Net models for comprehensive , 8-class land cover mapping, covering more than 123 billion pixels over the state of Mississippi, USA. Detailed qualitative and quantitative analysis revealed accurate mapping of open water and forested areas, while highlighting challenges in accurate delineation between cropland, herbaceous, and barren land cover types. These results show that self-supervised learning is an effective strategy for reducing the need for large volumes of manually annotated data, directly addressing a major limitation to high spatial resolution land cover mapping at scale.
@@ 1 Introduction
Land cover data products provide valuable information about the characteristics of the Earth’s surface for a wide variety of applications, including agriculture, urban planning, and environmental monitoring [107, 121, 136, 4]. Research into automated methods for land cover classification from remotely sensed data has been ongoing for several decades, with a majority of methods relying on supervised learning techniques to classify land cover using multispectral aerial or satellite imagery [13, 12, 35, 156]. Many early efforts into automated land cover classification relied on parametric models to classify individual pixels in coarse resolution satellite imagery based on their spectral signatures or temporal trends [26, 57, 120, 83]. Pixel-wise classification approaches matured in the late 1990s and 2000s as computational resources became more widely available and more capable non-parametric machine learning models were adopted for such tasks [78, 11, 97, 50, 34]. However, despite the increased predictive capacity of these non-parametric models, classification of very high spatial resolution (VHSR, ¡) remained challenging, primarily due to the inherent complexity of land cover patterns at these scales and the increased spectral variability within classes [113, 38]. Woodcock and Strahler [134] demonstrated that local spectral variability in land cover classes substantially increases as the spatial resolution of imagery increases beyond , increasing the inherent intra-class variance and making pixel-wise classification of VHSR imagery challenging [22]. This is problematic, as VHSR land cover products offer detailed information about the Earth’s surface with high spatial accuracy, which is valuable for applications that require fine-grained information about land cover types and their spatial distribution [67, 33, 41].
Object-based image analysis (OBIA) paradigms emerged in the early 2000s in an effort to address the shortcomings of pixel-wise methods by aggregating neighboring pixels with similar characteristics into objects using a segmentation algorithm; instead of classifying individual pixels, the classification algorithm operates on the objects [7, 8]. By incorporating spatial context into the classification process, OBIA land cover classification algorithms can outperform their pixel-based counterparts at high spatial resolutions [99, 3, 117, 60]. However, the performance of OBIA classifiers heavily relies on the segmentation algorithm’s ability to meaningfully delineate objects in the source imagery in an accurate and reliable fashion. Selecting and calibrating a segmentation algorithm to robustly segment objects is an ill-defined process that requires substantial human supervision before a supervised classifier can be trained [81]. A poor choice of segmentation algorithm or its parameters can substantially degrade performance, as objects extracted by the segmentation algorithm may group multiple objects with different land cover classes together or may fragment a single object into multiple segments [48]. While multi-scale segmentation algorithms attempt to address these shortcomings, they introduce additional complexity into the classification process that makes developing and scaling these models for operational land cover mapping difficult [29, 90, 39].
Subsequently, the introduction of deep learning-driven convolutional neural networks (CNNs) for general-purpose image processing tasks in the broader computer vision research community sparked substantial interest from researchers in remote sensing looking to consolidate the multi-stage classification process of OBIA techniques into an end-to-end modeling framework capable of incorporating both spectral and spatial features into predictions. CNNs utilize stacked convolutional kernels with learnable parameters to extract spatially-organized feature maps corresponding to the presence of spectral and spatial features in the input [62]. Originally designed for image classification, Long et al. [76] demonstrated that CNNs could generate dense pixel-wise classification maps by upsampling these feature maps to match the resolution of the input, then applying a simple linear classifier to each pixel in the upsampled feature maps. Despite the relative simplicity of this approach compared to existing methods, these fully convolutional networks (FCNs) achieved high performance on difficult general-purpose semantic segmentation tasks. The design of FCNs became a starting point for research into deep learning models designed explicitly for semantic segmentation, and works centered around integrating these models for classification of remotely sensed data soon followed [148]. These semantic segmentation networks have quickly become the state-of-the-art for VHSR land cover classification [109, 27, 130, 69, 150].
@@
Nevertheless, these models require far more labeled training data compared to traditional approaches, serving as a significant barrier to their widespread adoption for operational land cover mapping at VHSR [151, 49]. This need for large amounts of ground-truth data for accurate training of deep learning models has led to extensive research into methods for reducing the amount of data necessary to train these model [137, 122, 85]. Most of these methods leverage transfer learning procedures, where the parameters of a model pre-trained using a large dataset are used to initialize a model that is subsequently fine-tuned on a smaller dataset that is more relevant to the target task [132, 157]. For example, a common approach to transfer learning for land cover classification tasks is to pre-train a model on a large general-purpose image classification dataset (e.g., ImageNet [106]) and then fine-tune the model on a smaller dataset of interest [98, 84, 94, 158]. Despite being effective, the domain shift between the source and target datasets can hinder the model’s performance, and the extent of this domain shift between general-purpose image classification datasets and VHSR remote sensing datasets is not well understood [82]. As fully-supervised pre-training requires large-scale labeled datasets, which are relatively scarce in the remote sensing domain for VHSR imagery, there is a need for alternative pre-training strategies that reduce the reliance on labeled data [127, 143].
@@
More recently, self-supervised pre-training methods have been proposed as an alternative to traditional fully-supervised pre-training methods, enabling models to learn useful spatial and spectral information from input imagery through the use of pretext tasks that do not require labeled data [135, 119, 18]. These techniques are particularly attractive for remote sensing applications, as large amounts of labeled imagery datasets are often unavailable, but source imagery is widely available due to the growing number of Earth observation satellites and aerial imagery platforms [79]. However, few studies have rigorously investigated the applicability of these methods to operational land cover mapping at VHSR, especially in scenarios where ground-truth data are scarce. Most research evaluating self-supervised models pre-trained with remote sensing data uses benchmark VHSR land cover datasets as a proxy to evaluate the quality of pre-trained models, with the ISPRS Potsdam [53] and Vaihingen [54] datasets seeing wide use for this purpose [115, 128, 37, 126]. However, these datasets contain relatively large amounts of labeled training data, and thus may not necessarily be representative of real-world scenarios where no available VHSR land cover data exists and costly manual annotation is required to create such data (e.g., by experience, a single 256256 image patch can take upwards of to label well). Furthermore, while there is growing momentum in the development of large remote sensing foundation models using self-supervised techniques [79, 138], there is still a critical need to systematically assess how these approaches can enhance VHSR land cover mapping performance in practical contexts, a gap this study directly addresses.
@@
Self-supervised approaches inspired by masked language modeling in natural language processing have been shown to be strong approaches for learning useful feature representations from images without the need for labeled data [42, 141]. Masked image modeling (MIM) methods take advantage of the Vision Transformer (ViT) patching mechanism to randomly mask out a portion of the input image and then train the model to reconstruct the missing patches from the visible patches [70, 131]. MIM techniques have become the predominant self-supervised pre-training paradigm for creating remote sensing foundation models, which are large pre-trained models that can be adapted for a wide variety of tasks and domains that ideally outperform models trained from scratch or use more traditional transfer learning approaches [126, 47]. However, while powerful, ViTs lack the inductive biases towards imagery that make CNNs effective at extracting spatial features from images [1]. This is problematic, as ViTs typically only begin to outperform CNNs when pre-trained on very large datasets [28], a practical consideration for researchers and practitioners who must balance both computational resources and data availability when developing models. On the other hand, contrastive pretext tasks train a model to maximize the similarity of embeddings produced by positive pairs (i.e., two randomly augmented views of the same image) while simultaneously minimizing the distance between negative pairs (i.e., two views from separate images). While contrastive methods are effective and are not restricted to ViTs like their MIM counterparts, they typically rely on large in-memory batch sizes to effectively learn useful representations due to the restrictions of contrastive loss functions. For example, ablation studies of the popular SimCLR method demonstrate that a batch size of 1024 or greater is necessary to achieve adequate performance [18]. Subsequently, using traditional contrastive pre-training requires large amounts of computational resources, in particular VRAM, for pre-training to be effective [15]. Momentum contrast (MoCo) is a variant of contrastive pre-training that uses a queue of embeddings to serve as negative pairs: instead of requiring all negative samples to be held in VRAM, a separate momentum encoder network is used to generate these embeddings that are then used in the contrastive loss. Crucially, this momentum encoder is continuously updated as a slow-moving average of the network being trained. This framework enables more compute-efficient contrastive pre-training [43, 20].
Further, self-distillation techniques, such as Bootstrap Your Own Latent (BYOL) [36] and DINO [14], are alternatives to MIM and contrastive pre-training that can be used to learn robust and transferable feature representations without the need for negative samples, large batch sizes, or ViT backbones, which are hard requirements for many other self-supervised learning methods [18, 15]. Such approaches largely revolve around maximizing the similarity between feature representations of different augmented views of the same input using a teacher-student framework [45]. While the student network is trained directly using standard backpropagation and gradient descent algorithms, the teacher’s parameters are updated using a momentum-based rule derived from the student’s parameters. Self-distillation methods are attractive for remote sensing applications due to their inherent flexibility and ability to learn strong feature representations using a variety of backbone architectures, including CNNs [21]. Given the architectural constraints of MIM, contrastive learning and self-distillation techniques offer compelling, architecture-agnostic solutions for pre-training CNNs on domain-specific, moderate-sized datasets [55].
@@
In this study, we developed an innovative label-efficient framework that addresses the challenge of operational land cover mapping at resolution under label-scarce scenarios through self-supervised learning. Specifically, we evaluated the use of BYOL and MoCoV2 as pretext tasks for self-supervised learning during pre-training, enabling a ResNet-101 convolutional encoder network [44] to learn to extract deep features from 377,921 unlabeled 256256 patches of color-infrared (CIR) aerial imagery without the need for full supervision using labeled data. We then transferred the learned encoder weights into multiple semantic segmentation architectures (FCN, U-Net, Attention U-Net, DeepLabV3+, UPerNet, PAN) that were subsequently fine-tuned using very small training dataset sizes (250, 500, and 750 256256 patches) using cross-validation. Evaluating the performance of these models showed that initializing a model with out-of-domain ImageNet weights then applying BYOL pre-training on the in-domain imagery led to increases in accuracy across all other pre-training protocols tested. We then classified 8 classes of land cover (open water, impervious structures, impervious surfaces, barren land, forest/woody vegetation, herbaceous/low vegetation, cultivated crops, and unclassified (i.e., shadow)) over the entire state of Mississippi, USA using an ensemble of the best-performing models, totaling over 123 billion pixels. Validation and accuracy assessment were performed using a spatially independent test set of 25,000 randomly sampled points with manually verified labels to ensure robust and unbiased evaluation of our approach and the resulting land cover map. Finally, we evaluated the generalization capacity of this self-supervised model ensemble by applying it to 2016 imagery with different characteristics and validating the resulting predicted land cover data against a separate ground truth dataset to show that the model can accurately classify imagery outside of the training dataset distribution. We demonstrated that these approaches can be used to train deep semantic segmentation models that are accurate, generalizable, and capable of yielding reasonable performance even under extreme low-data scenarios, even when only 3 bands (red, green, and near-infrared) are available in the input imagery. Our ultimate aim is to reduce practical barriers to producing high spatial resolution land cover maps while contributing to ongoing efforts to evaluate and contextualize the implications of self-supervised learning techniques for remote sensing purposes.
@@ 2.1 NAIP aerial imagery
The USDA’s National Agricultural Imagery Program (NAIP) is a U.S. federal project that provides a publicly available archive of high-resolution multispectral imagery over the conterminous United States, with a spatial resolution of , a four-band spectral resolution (near-infrared, red, green, and blue), delivered as 8-bit orthorectified tiles [32]. NAIP imagery is collected on a state-by-state basis during the agricultural growing season, and each state is imaged once every 2–3 years. The lack of use restrictions on NAIP imagery has made it a popular choice for high spatial resolution remote sensing applications in the United States, including building footprint extraction [51], forest measurements [46, 25], and land cover classification [88, 87, 40, 86]. NAIP imagery is provided as projected rasters according to each raster’s appropriate UTM zone, but Mississippi spans across 2 UTM zones (15N/16N); for consistency across the dataset, we reprojected and resampled all imagery to the Mississippi Transverse Mercator coordinate reference system (EPSG:3813) at resolution using bilinear interpolation. CIR composites for years 2016 and 2023 were generated using the near-infrared, red, and green bands to emphasize vegetation features in the imagery. We selected three-band images for compatibility with BYOL and MoCoV2, which are designed for three-channel inputs, and for easy comparison with models using ImageNet pre-trained weights, which also expect three-channel inputs. In total, of NAIP composite imagery across 2016 and 2023 was downloaded and processed for this study. A visualization of the 2023 NAIP imagery used over the study area is provided in Figure 1.
@@ 2.2 NLCD land cover
The stratification approach proposed in this work (Section 3.1) relies on an existing land cover product to select strata from which training samples are drawn. For this purpose, we used the USGS’s National Land Cover Database (NLCD) to provide the necessary land cover information for stratification. The NLCD has historically been among the most widely used land cover datasets in the United States, supporting numerous applications in human and ecosystem health studies, urban planning, natural hazard assessment, hydrology, and land management [110]. The latest version of the NLCD (Annual NLCD) provides a yearly 20-class land cover product primarily derived from Landsat imagery at resolution from 1985 to 2024 [123]. While the spatial resolution of the NLCD land cover product is substantially coarser than the land cover product we produced in this study, the NLCD provides a comprehensive land cover product at a national scale and has a history of consistent use in the literature [144, 6, 92], making it a suitable choice to provide land cover information to guide the stratification process. To align with the collection year of the NAIP imagery used to train the model, we used 2023 data over the state of Mississippi.